# A Reduced Basis Decomposition Approach to Efficient Data Collection in Pairwise Comparison Studies

### v1

@@ Abstract
Comparative judgement studies elicit quality assessments through pairwise comparisons, typically analysed using the Bradley-Terry model. A challenge in these studies is experimental design, specifically, determining the optimal pairs to compare to maximize statistical efficiency. Constructing static experimental designs for these studies requires spectral decomposition of a covariance matrix over pairs of pairs, which becomes computationally infeasible for studies with more than approximately 150 objects. We propose a scalable method based on reduced basis decomposition that bypasses explicit construction of this matrix, achieving computational savings of two to three orders of magnitude. We establish eigenvalue bounds guaranteeing approximation quality and characterise the rank structure of the design matrix. Simulations demonstrate speedup factors exceeding 100 for studies with 64 or more objects, with negligible approximation error. We apply the method to construct designs for a 452-region spatial study in under 7 minutes and enable real-time design updates for classroom peer assessment, reducing computation time from 15 minutes to 15 seconds.
keywords:
Comparative judgement , Bradley–Terry , Matrix Approximation , Scalable algorithms
[…]
organization=School of Mathematics,addressline=University of Birmingham, city=Edgbaston, postcode=B15 2TT, country=United Kingdom
@@ 1 Introduction
Comparative judgement is a type of study that has been widely used across the social sciences and humanities. In its simplest form, study participants are shown objects and asked which object has a higher quality. From these pairwise comparisons, a statistical model can be used to estimate the quality of each object and construct a ranked ordering of the objects. For example, in Education research, comparative judgement studies have been carried out with teachers comparing students’ mathematics homework based on which showed a better understanding of fractions (Jones2015); in Linguistics, language learners have been asked which of a pair of words in a foreign language was more difficult to learn (Bisson2022); and in one Philosophy study, participants were shown pairs of vehicles and asked which was worse a violation of a sign that reads “no vehicles in the park”, in order to understand rule-violations (Tanswell23).
@@
The Bradley–Terry model (Brad52) is perhaps the most commonly used statistical model to estimate the qualities of the objects in a comparative judgement study. This model assigns a quality parameter to each object in the study and estimates the value of these parameters from the pairwise comparisons. One limitation of the Bradley–Terry model is the difficulty of generating an experimental design. The main experimental design element in a comparative judgement study is deciding which pairs of objects to ask participants about. Experiential designs can take the form of scheduling distributions, which are a distribution over all possible pairs of objects, and determine the probability each pair of objects is shown to the study participant. Constructing the scheduling distribution for comparative judgement studies requires us to model the dependencies between comparisons, necessitating the consideration of not just pairs of objects, but pairs of pairs of objects. This means the associated matrices are prohibitively large, dense, and exhibit a non-standard structure. Indeed, many studies are simply too large for current experimental design methods. A recent meta-analysis of 41 comparative judgement studies reported that 40% contained more than 100 objects, while some citizen science studies have involved several thousand objects (Kinnear2025).
@@
Well-designed experiments can substantially reduce the amount of data required, making comparative judgement studies more feasible in practice. Experimental designs for the Bradley–Terry models broadly fall into two categories: online designs (known as adaptive comparative judgement) and static designs that are constructed before the study begins. Pollitt2012 created an online experimental design, known as adaptive comparative judgement, where the objects featured in upcoming comparisons are determined based on comparisons made earlier in the study. Adaptive designs have come under criticism for two reasons. The first is that comparisons towards the end of the study are typically more difficult to make, as the adaptive design prioritises pairs that are close in quality (Jones2023). This can introduce an unwanted study effect, where data collected later on in the study period requires more effort to collect than data earlier on in the study. The second is that adaptive comparative judgement artificially inflates the reliability of the resulting parameters estimates (Bramley2018). Additionally, a limitation of adaptive designs is that they are only based on the quality parameters, and do not take into account any prior information or covariates.
@@
We focus on studies where the design is static and must be pre-calculated to account for specific constraints. This includes scenarios involving prior information, such as spatial correlations between objects, as well as sequential designs where the results of one round inform the scheduling of the next. Example of multi-phase adaptive designs include sporting contests and educational peer-assessment activities, where students evaluate work in distinct stages separated by other learning activities.
@@
Several methods exist for creating static designs, but can only be used for small studies or where no prior information is taken into account. Grass08 were able to analytically derive a static experimental design, but this is limited to studies with at most three objects. More recently, Roettger2025 develop a method for constructing the D-optimal design, but not in a Bayesian framework. Seymour25 developed a static design for studies where the objects are a priori correlated. This method creates a scheduling distribution that specifies the probability of each pair of objects being compared. However, constructing this distribution is computationally expensive and can only be computed for studies with fewer than 150 objects, and it may take several hours or days to construct the distribution. These limitations comes about because we need to construct and decompose the large, dense matrix describing the correlation between pairs of pairs of objects.
@@ 2.1 Likelihood function and inference
By Bayes’ theorem, the posterior distribution is given by $\pi(\boldsymbol{\lambda}\mid\boldsymbol{y})\propto f(\boldsymbol{y}\mid\boldsymbol{\lambda})\pi(\boldsymbol{\lambda})$. To allow for prior correlation between the object quality parameters and to allow for an offline experimental design, we place a multivariate normal prior distribution on the quality parameters. This distribution is given by $\boldsymbol{\lambda}\sim\rm MVN(\boldsymbol{\mu},C)$. The structure of the covariance matrix $C$ is a modelling choice and should reflect a priori assumptions about how the qualities of the objects are related.
@@ 2.2 Experimental design
In this paper, we focus on the design proposed in Seymour25. This is a principal component based design, where the scheduling distribution places higher probability on pairs of objects that explain higher prior variance. This design has been shown to reduce the number of comparisons that need to be collected. Since the Bradley–Terry model relies on pairwise differences in quality parameters, we can derive the distribution of these differences. This is derived via an affine transformation of the prior distribution $\boldsymbol{\lambda}\sim\mathcal{N}(\boldsymbol{\mu},C)$, yielding:
@@
Here, the vector $\boldsymbol{\lambda}_{\text{diff}}\in\mathbb{R}^{N(N-1)/2}$ consists of the quality differences for each pair of objects, such that the entry corresponding to pair $(i,j)$ is $\lambda_{i}-\lambda_{j}$ (for $i<j$). The vector $\boldsymbol{\nu}$ represents the corresponding differences in means, and $\Delta\in\mathbb{R}^{\frac{N(N-1)}{2}\times\frac{N(N-1)}{2}}$ is the covariance matrix capturing the dependence structure between pairwise comparisons. The entries of $\Delta$, representing the covariance between two pairs $(i,j)$ and $(k,l)$, are given by:
@@
We perform a spectral decomposition of the covariance matrix such that $\Delta=U\Psi U^{T}$, where $\Psi$ is a diagonal matrix of eigenvalues and the columns of $U$ contain the corresponding eigenvectors. We order the eigenpairs $\{\psi^{(c)},\mathbf{u}^{(c)}\}$ in descending order such that $\psi^{(c)}\geq\psi^{(c+1)}\geq 0$, ensuring non-negativity as $\Delta$ is positive semidefinite. The eigenvectors $\{\mathbf{u}^{(c)}\}$ form an orthogonal basis for $\mathbb{R}^{M}$, where each vector represents a principal component (Mardia79).
@@
A workflow for the construction of $\mathcal{S}$ is shown in Algorithm 1.
@@ 3 Reduced basis decomposition for the Bradley–Terry model
Current static experimental designs become infeasible for large $N$ due to the quartic growth in memory and polynomial growth in compute time. Because $\Delta$ is a dense matrix of size $\frac{N(N-1)}{2}\times\frac{N(N-1)}{2}$, it imposes an $\mathcal{O}(N^{4})$ storage overhead. This volume renders standard PCA impractical with its $\mathcal{O}(N^{6})$ complexity (jolliffe2011principal). Furthermore, even truncated variants such as randomized PCA (eckart1936approximation; halko2011finding) and Lanczos algorithms (lehoucq1998arpack) fail to sufficiently alleviate the cost, requiring $\mathcal{O}(N^{4}d)$ operations. To address this, we propose a novel computational framework to efficiently construct $\Delta$ and compute its spectral decomposition by exploiting the reduced basis decomposition (RBD) method. Our approach has three main components:
@@ 3.2 Reduced-basis spectral decomposition of $\boldsymbol{E}$
The RBD method is based on a greedy algorithm commonly used in model-reduction techniques (ngoc2005certified; chen2013reduced; rozza2008reduced). It builds the reduced space dimension-by-dimension: At each step, the greedy algorithm identifies the column of $\boldsymbol{E}$ that is least well represented by the current reduced space and adds it as a new basis vector. This choice is guided by a computable error indicator that measures the discrepancy between each column of $\boldsymbol{E}$ and its projection onto the current reduced space. By repeatedly selecting the direction with the largest error, the algorithm ensures that the reduced basis improves systematically and targets the components of $\boldsymbol{E}$ that are most difficult to approximate.
@@ 3.4 Theoretical analysis
By eq. 12 and eq. 13, $\{\sigma_{i}\}_{i=1}^{d}$ are the eigenvalues of $\widetilde{\boldsymbol{C}}$ and the largest $d$ eigenvalues of $\widetilde{\Delta}$. By Cauchy’s interlacing theorem (bellman1997introduction; horn2012matrix; parlett1998symmetric),
@@ 4.1 Scalability in size of study
We simulate scenarios with increasing $N$, ranging from small-scale studies ($N=8$) to high-dimensional cases ($N=128$) where standard decomposition techniques typically become intractable. To investigate the use of the method for different kinds of prior covariance structures, we use simulate different covariance matrices. The first is a Graph Laplacian covariance structure, where we simulate an Erdős–Rényi graph with $N$ nodes and the probability of an edge being placed between two nodes $p=0.5$, i.e. $G(N,0.5)$. We used $p=0.5$ to ensure the resulting graphs are sufficiently connected, while retaining enough sparsity to be non-trivial. The Graph Laplacian matrix is given by $C=(D-A+I)^{-1}$, where $A$ is the adjacency matrix. This enforced smoothness between the study objects and can be used in Gaussian Markov Random Fields. The second is a Toeplitz covariance matrix, where $C_{ij}=0.5^{|i-j|}$, which can be used where there is a prior autoregressive or ordering assumption. The final covariance structure is a Inverse-Wishart distribution with the identity matrix $I$ as the scale matrix, which represents studies where there is weakly informative prior information. For structure, we normalise the resulting matrix by $C^{\prime}=D^{-\frac{1}{2}}CD^{-\frac{1}{2}}$, where $D$ is a diagonal matrix containing the elements on the diagonal of $C$. This allows us to provide a more fair comparison between simulations, as the variance of each objects’ quality is the same.
@@
Figure 2: Comparison of KL diveregnce for studies with different numbers of objects (N) in. The subfigures show (a) Laplacian, (b) Toeplitz, and (c) Inverse Wishart. Note the y-axes are on the log scale.
@@ 4.2 Sensitivity to the sparsity of the prior covariance matrix
In Table 2, we report the mean computation time (in seconds), mean speedup (the computation time for standard method divded by the computation time for the RBD) and the mean Kullback-Leibler (KL) divergence for each method. For all values of sizes and sparsities of the covariance matrix, we see the KL divergence is negligible and at most $1.11\times 10^{-5}$. This suggests our method is not sensitive to the sparsity of the covariance matrix, and is suitable for complex prior distributions. We also see the time taken to compute the scheduling distribution does not vary as the level of sparsity changes. For the studies with 128 objects, the RBD method takes 2-3 seconds for all values of $p$, compared the the standard methods which takes 5-8 minutes.
@@ 5.1 Scheduling for large spatial studies
To demonstrate the utility of the proposed Reduced Basis Decomposition method in a real-world setting, we consider the large-scale comparative judgement study presented by Seymour2022. The objective of the study was to use comparative judgement to quantify levels of urban deprivation across Dar es Salaam, Tanzania, overcoming the limitations of traditional household surveys which are often costly, slow, and prone to rapid obsolescence in fast-growing cities.
The study employed a pairwise comparison approach to elicit local knowledge from citizens. The objects of interest were $N=452$ administrative sub-wards within the city. Judges were presented with pairs of sub-wards (displayed via imagery or identified by name) and asked to select the more affluent of the two. Seymour2022 collected approximately 75,000 comparisons to achieve stable estimates, utilizing a spatial prior distribution to borrow strength between neighbouring regions. A scheduling algorithm could have reduced the number of comparisons that were collected, but it was not possible to construct without our RBD method, especially in the field in Tanzania.
@@
We follow Seymour2022 and construct an adjacency matrix $A$ from the sub-wards of Dar es Salaam by treating the wards as nodes and placing edges between geographically adjacent wards. The prior distribution on the deprivation parameters of each ward is given by $\boldsymbol{\lambda}\sim N(\boldsymbol{0},\,C)$, where $C=D^{-\frac{1}{2}}\Lambda D^{-\frac{1}{2}}$, $\Lambda=e^{A}$, $A$ is the network’s adjacency matrix, and $D$ is a diagonal matrix containing the elements on the diagonal of $\Lambda$. Using the matrix exponential of the adjacency matrix as the basis of the covariance matrix assigns higher prior covariance to pairs of subwards that are well-connected in the network, and low prior covariance to pairs that are poorly connected.

### v2

@@ Abstract
Comparative judgement studies elicit quality assessments of objects through pairwise comparisons, typically analysed using the Bradley-Terry model. A challenge in these studies is experimental design, specifically, determining the optimal pairs to compare to maximize statistical efficiency. Constructing static experimental designs for these studies requires spectral decomposition of a covariance matrix over pairs of pairs, which becomes computationally infeasible for studies with a large number of objects. We propose a scalable method based on reduced basis decomposition that bypasses explicit construction of this matrix, achieving computational savings of two to three orders of magnitude. We establish eigenvalue bounds guaranteeing approximation quality and characterise the rank structure of the design matrix. Simulations demonstrate speedup factors exceeding 100 for studies with 64 or more objects, with negligible approximation error. We apply the method to construct designs for a 452-region spatial study in under 7 minutes, which was not previously possible, and enable real-time design updates for classroom peer assessment, reducing computation time from 15 minutes to 15 seconds.
@@ 1 Introduction
Comparative judgement is a type of study that has been widely used across the social sciences and humanities. In its simplest form, study participants are shown objects and asked which object has a higher quality. From these pairwise comparisons, a statistical model can be used to estimate the quality of each object and construct a ranked ordering of the objects. For example, in education research, comparative judgement studies have been carried out with teachers comparing students’ mathematics homework based on which showed a better understanding of fractions (Jones and Wheadon, 2015); in Linguistics, language learners have been asked which of a pair of words in a foreign language was more difficult to learn (Bisson, 2022); and in one Philosophy study, participants were shown pairs of vehicles and asked which was worse a violation of a sign that reads “no vehicles in the park”, in order to understand rule-violations (Tanswell et al., 2023).
@@
The Bradley–Terry model (Bradley and Terry, 1952) is perhaps the most commonly used statistical model to estimate the qualities of the objects in a comparative judgement study. This model assigns a quality parameter to each object in the study and estimates the value of these parameters from the pairwise comparisons. One limitation of the Bradley–Terry model is the difficulty of generating an experimental design. The main experimental design element in a comparative judgement study is deciding which pairs of objects to ask participants about. Experimental designs can take the form of scheduling distributions, which are a distribution over all possible pairs of objects, and determine the probability each pair of objects is shown to the study participant. Constructing the scheduling distribution for comparative judgement studies requires us to model the dependencies between comparisons, necessitating the consideration of not just pairs of objects, but pairs of pairs of objects. This means the associated matrices are prohibitively large, dense, and exhibit a non-standard structure. Indeed, many studies are simply too large for current experimental design methods. A recent meta-analysis of 41 comparative judgement studies reported that 40% contained more than 100 objects, while some citizen science studies have involved several thousand objects (Kinnear et al., 2025).
@@
Well-designed experiments can substantially reduce the amount of data required, making comparative judgement studies more feasible in practice. Experimental designs for the Bradley–Terry models broadly fall into two categories: online designs (known as adaptive comparative judgement) and static designs that are constructed before the study begins. Pollitt (2012) created an online experimental design, known as adaptive comparative judgement, where the objects featured in upcoming comparisons are determined based on comparisons made earlier in the study. Adaptive designs have come under criticism for two reasons. The first is that comparisons towards the end of the study are typically more difficult to make, as the adaptive design prioritises pairs that are close in quality (Jones and Davies, 2023). This can introduce an unwanted study effect, where data collected later on in the study period requires more effort to collect than data earlier on in the study. The second is that adaptive comparative judgement artificially inflates the reliability of the resulting parameter estimates (Bramley and Vitello, 2018). Additionally, a limitation of adaptive designs is that they are only based on the quality parameters, and do not take into account any prior information or covariates.
@@
We focus on studies where the design is static and must be pre-calculated to account for specific constraints. This includes scenarios involving prior information, such as spatial correlations between objects, as well as sequential designs where the results of one round inform the scheduling of the next. Examples of sequential designs include sporting contests and educational peer-assessment activities, where students evaluate work in distinct stages separated by other learning activities.
@@
Several methods exist for creating static designs, but can only be used for small studies or where no prior information is taken into account. Grasshoff and Schwabe (2008) were able to analytically derive a static experimental design, but this is limited to studies with at most three objects. More recently, Röttger et al. (2025) develop a method for constructing the D-optimal design, but not in a Bayesian framework. Seymour et al. (2025) developed a static design for studies where the objects are a priori correlated. This method creates a scheduling distribution that specifies the probability of each pair of objects being compared. However, constructing this distribution is computationally expensive and can only be computed for studies with fewer than 150 objects, and it may take several hours or days to construct the distribution. These limitations come about because we need to construct and decompose the large, dense matrix describing the correlation between pairs of pairs of objects.
@@ 2.1 Likelihood function and inference
By Bayes’ theorem, the posterior distribution is given by $\pi(\boldsymbol{\lambda}\mid\boldsymbol{y})\propto f(\boldsymbol{y}\mid\boldsymbol{\lambda})\pi(\boldsymbol{\lambda})$. To allow for prior correlation between the object quality parameters and to allow for an offline experimental design, we place a multivariate normal prior distribution on the quality parameters. This distribution is given by $\boldsymbol{\lambda}\sim\rm MVN(\boldsymbol{\mu},C)$.
The structure of the covariance matrix $C$ is a modelling choice and should reflect a priori assumptions about how the qualities of the objects are related. In some settings where data is collected in rounds, for example student’s abilities in educational assessments or sporting teams quality in tournaments, the prior covariance matrix can be constructed from the results of previous rounds (e.g. via the Hessian of the maximum likelihood estimates, or the posterior covariance matrix). In contexts where there study items have covariate information, the covariates can be form the prior covariance structure, particularly if these are spatial covariates. Finally, in contexts where there is a natural prior ranking to the objects, this can be included in a Toeplitz type structure, where the covariance between items $i$ and $j$ depend on |ij||i-j|.
@@ 2.2 Experimental design
In this paper, we focus on the design proposed in Seymour et al. (2025). This is a principal component based design, where the scheduling distribution places higher probability on pairs of objects that explain higher prior variance. This design has been shown to reduce the number of comparisons that need to be collected. Since the Bradley–Terry model relies on pairwise differences in quality parameters, we can derive the distribution of these differences. This is derived via an affine transformation of the prior distribution $\boldsymbol{\lambda}\sim\mathcal{N}(\boldsymbol{\mu},C)$, yielding:
@@
Here, the vector $\boldsymbol{\lambda}_{\text{diff}}\in\mathbb{R}^{N(N-1)/2}$ consists of the quality differences for each pair of objects, such that the entry corresponding to pair $(i,j)$ is $\lambda_{i}-\lambda_{j}$ for $i<j$, where the condition $i<j$ is used solely to index unique pairs and does not imply an ordering of the quality parameters. The vector $\boldsymbol{\nu}$ represents the corresponding differences in means, and $\Delta\in\mathbb{R}^{\frac{N(N-1)}{2}\times\frac{N(N-1)}{2}}$ is the covariance matrix capturing the dependence structure between pairwise comparisons. The entries of $\Delta$, representing the covariance between two pairs $(i,j)$ and $(k,l)$, are given by:
@@
We perform a spectral decomposition of the covariance matrix such that $\Delta=U\Psi U^{T}$, where $\Psi$ is a diagonal matrix of eigenvalues and the columns of $U$ contain the corresponding eigenvectors. We order the eigenpairs $\{\psi^{(c)},\mathbf{u}^{(c)}\}$ in descending order such that $\psi^{(c)}\geq\psi^{(c+1)}\geq 0$, ensuring non-negativity as $\Delta$ is positive semidefinite. The eigenvectors $\{\mathbf{u}^{(c)}\}$ form an orthogonal basis for $\mathbb{R}^{M}$, where each vector represents a principal component (Mardia et al., 1979).
@@
A workflow for the construction of $\mathcal{S}$ is shown in Algorithm 1. We note that this Algorithm only provides information about which comparisons should be collected and not how many comparisons need to be collected. This is still an area of active research in comparative judgement, however Kinnear et al. (2025) suggest that collecting ten times as many comparisons as there are items is required for reliable estimates for the quality parameters.
@@ 3 Reduced basis decomposition for the Bradley–Terry model
Current static experimental designs become infeasible for large $N$ due to the quartic growth in memory and polynomial growth in compute time. Because $\Delta$ is a dense matrix of size $\frac{N(N-1)}{2}\times\frac{N(N-1)}{2}$, it imposes an $\mathcal{O}(N^{4})$ storage overhead. This volume renders standard PCA impractical with its $\mathcal{O}(N^{6})$ complexity (Jolliffe, 2011). Furthermore, even truncated variants such as randomized PCA (Eckart and Young, 1936; Halko et al., 2011) and Lanczos algorithms (Lehoucq et al., 1998) fail to sufficiently alleviate the cost, requiring $\mathcal{O}(N^{4}d)$ operations. To address this, we propose a novel computational framework to efficiently construct $\Delta$ and compute its spectral decomposition by exploiting the reduced basis decomposition (RBD) method. Our approach has three main components:
@@ 3.2 Reduced-basis spectral decomposition of $\boldsymbol{E}$
The RBD method is based on a greedy algorithm commonly used in model-reduction techniques (Ngoc Cuong et al., 2005; Chen and Gottlieb, 2013; Rozza et al., 2008). It builds the reduced space dimension-by-dimension: At each step, the greedy algorithm identifies the column of $\boldsymbol{E}$ that is least well represented by the current reduced space and adds it as a new basis vector. This choice is guided by a computable error indicator that measures the discrepancy between each column of $\boldsymbol{E}$ and its projection onto the current reduced space. By repeatedly selecting the direction with the largest error, the algorithm ensures that the reduced basis improves systematically and targets the components of $\boldsymbol{E}$ that are most difficult to approximate.
@@ 3.4 Theoretical analysis
By eq. 12 and eq. 13, $\{\sigma_{i}\}_{i=1}^{d}$ are the eigenvalues of $\widetilde{\boldsymbol{C}}$ and the largest $d$ eigenvalues of $\widetilde{\Delta}$. By Cauchy’s interlacing theorem (Bellman, 1997; Horn and Johnson, 2012; Parlett, 1998),
@@ 4.1 Scalability in size of study
We simulate scenarios with increasing $N$, ranging from small-scale studies ($N=8$) to high-dimensional cases ($N=128$) where standard decomposition techniques typically become intractable. To investigate the use of the method for different kinds of prior covariance structures, we simulate different covariance matrices. The first is a Graph Laplacian covariance structure, where we simulate an Erdős–Rényi graph with $N$ nodes and the probability of an edge being placed between two nodes $p=0.5$, i.e. $G(N,0.5)$. We used $p=0.5$ to ensure the resulting graphs are sufficiently connected, while retaining enough sparsity to be non-trivial. The Graph Laplacian matrix is given by $C=(D-A+I)^{-1}$, where $A$ is the adjacency matrix. This enforced smoothness between the study objects and can be used in Gaussian Markov Random Fields. The second is a Toeplitz covariance matrix, where $C_{ij}=0.5^{|i-j|}$, which can be used where there is a prior autoregressive or ordering assumption. The final covariance structure is a Inverse-Wishart distribution with the identity matrix $I$ as the scale matrix, which represents studies where there is weakly informative prior information. For structure, we normalise the resulting matrix by $C^{\prime}=D^{-\frac{1}{2}}CD^{-\frac{1}{2}}$, where $D$ is a diagonal matrix containing the elements on the diagonal of $C$. This allows us to provide a more fair comparison between simulations, as the variance of each objects’ quality is the same.
@@
Figure 2: Comparison of KL diveregnce for studies with different numbers of objects (N). The subfigures show (a) Laplacian, (b) Toeplitz, and (c) Inverse Wishart. Note the y-axes are on the log scale.
@@ 4.2 Sensitivity to the sparsity of the prior covariance matrix
In Table 2, we report the mean computation time (in seconds), mean speedup (the computation time for standard method divided by the computation time for the RBD) and the mean Kullback-Leibler (KL) divergence for each method. For all values of sizes and sparsities of the covariance matrix, we see the KL divergence is negligible and at most $1.11\times 10^{-5}$. This suggests our method is not sensitive to the sparsity of the covariance matrix, and is suitable for complex prior distributions. We also see the time taken to compute the scheduling distribution does not vary as the level of sparsity changes. For the studies with 128 objects, the RBD method takes 2-3 seconds for all values of $p$, compared to the standard methods which takes 5-8 minutes.
@@ 5.1 Scheduling for large spatial studies
To demonstrate the utility of the proposed Reduced Basis Decomposition method in a real-world setting, we consider the large-scale comparative judgement study presented by Seymour et al. (2022). The objective of the study was to use comparative judgement to quantify levels of urban deprivation across Dar es Salaam, Tanzania, overcoming the limitations of traditional household surveys which are often costly, slow, and prone to rapid obsolescence in fast-growing cities. In the study, spatial information was used to construct the prior covariance matrix.
The study employed a pairwise comparison approach to elicit local knowledge from citizens. The objects of interest were $N=452$ administrative sub-wards within the city. Judges were presented with pairs of sub-wards (displayed via imagery or identified by name) and asked to select the more affluent of the two. Seymour et al. (2022) collected approximately 75,000 comparisons to achieve stable estimates, utilizing a spatial prior distribution to borrow strength between neighbouring regions. A scheduling algorithm could have reduced the number of comparisons that were collected, but it was not possible to construct without our RBD method, especially in the field in Tanzania.
@@
We follow Seymour et al. (2022) and construct an adjacency matrix $A$ from the sub-wards of Dar es Salaam by treating the wards as nodes and placing edges between geographically adjacent wards. The prior distribution on the deprivation parameters of each ward is given by $\boldsymbol{\lambda}\sim N(\boldsymbol{0},\,C)$, where $C=D^{-\frac{1}{2}}\Lambda D^{-\frac{1}{2}}$, $\Lambda=e^{A}$, $A$ is the network’s adjacency matrix, and $D$ is a diagonal matrix containing the elements on the diagonal of $\Lambda$. Using the matrix exponential of the adjacency matrix as the basis of the covariance matrix assigns higher prior covariance to pairs of subwards that are well-connected in the network, and low prior covariance to pairs that are poorly connected.